OpenStudy (anonymous):

Braden improves his score with each practice quiz he takes. The number of incorrect answers can be represented by the expression 20 – 5n where n represents the number of practice quizzes he has taken. How many answers will he get incorrect after taking 2 practice quizzes? 30 13 10 5

OpenStudy (anonymous):

He will get 5 * 2 = 10 answers wrong.

OpenStudy (anonymous):

so is dhat the rite answer

OpenStudy (anonymous):

10 is the right answer as 20 - 5 * 2 = 20 - 10 = 10

OpenStudy (anonymous):

okay thankz